OpenStudy (anonymous):

The sum of three numbers is 62. The second number is equal to the first number diminished by 4. The third number is four times the first. What are the numbers? Which three expressions could be used to represent the numbers? n; n - 4; 4n n; 4 - n; n + 4 n; n/4; n + 4

OpenStudy (anonymous):

n, n-4, 4n so n+(n-4) +4n=62 so, 2n-4+4n=62 so,6n-4=62 so, 6n=66 so6n/6=66/6 n=11 SO, the number are 11, 7 and 44 which add up to 62

OpenStudy (anonymous):

you ok Corey?

OpenStudy (anonymous):

ya